OpenStudy (anonymous):

is represented square root as @ then wouldnt -@100 equal both -10 and 10?

OpenStudy (anonymous):

because @100 equals -10 and 10 so if i multiplied both with - the awnsers would still be -10 and 10 just switched around respectively

OpenStudy (phi):

I would guess that if a sign is specified, they mean take the negative square root. example: \[ - \sqrt{4} = -2 \]

OpenStudy (phi):

It comes down to, what did the writer mean ?

OpenStudy (anonymous):

ok what if i did it like this -(\[\sqrt{100}\])

OpenStudy (anonymous):

i mean -(\[\sqrt{100}\])

OpenStudy (phi):

if they wanted you to consider both roots, it would be clearer to write \[ ± \sqrt{100} \] if they don't write that, but rather \[ - \sqrt{100} \] I would interpret the second way to mean "take only the negative root"

OpenStudy (anonymous):

ok thank you

OpenStudy (phi):

It helps to know the context of the expression. Whoever wrote this should be trying to be clear (as opposed to confusing).